# Contracting with a Mechanism Designer

### v1

@@ 6 Intermediary with Anonymous Pricing
Agents have identical expected reward profiles $\bm{\gamma}$ and contribution distributions $F$ and the mechanism space of the intermediary is restricted to anonymous pricing mechanisms.
@@
Under 6.1, we can obtain that given a contract $\bm{t}\in\mathbb{R}_{+}^{m}$, the intermediary’s expected utility under the anonymous price $p$ can be expressed as $\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\cdot\operatorname*{\mathrm{Pr}}\!\left[{\exists i\in[n],s.t.,c_{i}\leq p}\right]$. Note that the probability that there exists an agent with cost no more than the posted price $p$ is equivalent to the probability that there exists an agent with contribution no less than $\left\langle\bm{t},\bm{\gamma}\right\rangle-p$. Therefore, the intermediary’s expected utility of an anonymous price $p$ is equal to $\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\cdot\left(1-F^{n}_{\bm{t}}\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\right)$. By definition of reserve price, the intermediary’s optimal anonymous price is $\left\langle\bm{t},\bm{\gamma}\right\rangle-\theta^{(n)}_{\bm{t}}$, where $\theta^{(n)}_{\bm{t}}$ is the reserve price of the distribution $F^{n}_{\bm{t}}$. Thus, the principal’s utility under contract $\bm{t}$ is equal to
[…]
Similar to results in Section 4, we show that when agents are ex ante symmetric, the optimal contract is linear and its induced utility for the principal can be considered as a variant of virtual value pricing (Theorem 6.1). The proof is similar to the one for Theorem 4.1, and is deferred to Section C.4.
@@
Under 6.1, there exists an optimal contract that is linear. In particular, the optimal utility of the principal (denoted by $U_{\mathrm{posted}}$) is the optimal objective value of the optimization program defined as follows:
@@
Under 6.1, Theorem 6.1 implies that the principal’s problem of designing an optimal contract remains a virtual value pricing problem as established part (iii) of Theorem 4.1. The only difference is that the corresponding virtual value function becomes that of the distribution $F^{n}$. Intuitively, this virtual value pricing can be viewed as posting a price to a single virtual agent whose value is equal to the first-order statistic among that of the real agents. If she buys at that price, a payment equal to her virtual value will be received. Note that the principal’s utility in this case is no more than that in the case where there is no restriction on the intermediary’s mechanism space. A natural question is that is there any monotonicity in the principal’s utility regarding two intermediaries with an inclusion relationship between their mechanism spaces? We offer two negative examples for this question in Appendix B.
We next analyze the impact of the intermediary measured by PoA and PoDM as we have done in Section 5.151515Here we define PoA and PoDM under 6.1 by $\mathrm{PoA}\triangleq\frac{U_{\textsc{fb}}}{U_{\mathrm{posted}}}$ and $\mathrm{PoDM}\triangleq\frac{U_{\textsc{sb}}}{U_{\mathrm{posted}}}$, where benchmarks $U_{\textsc{fb}},U_{\textsc{sb}}$ follow the same definitions in Section 5. Interestingly, as shown in Theorems 6.2 and 6.3, even when the intermediary is restricted to the space of anonymous pricing mechanisms, the (worst-case) PoDM and PoA are the same as the setting where the intermediary can design any DSIC, ex-post IR mechanisms. The proofs of both theorems are similar to their analogy in Section 5 and thus are deferred into Sections C.5 and C.6, respectively.
@@
Under 6.1, if contribution distributions $F$ are regular with support $\mathtt{supp}(F)=\left[\underline{w},\bar{w}\right]$ $\left(0<\underline{w}\leq\bar{w}\leq r\right)$, if $1\leq\bar{w}/\underline{w}\leq\kappa$, the tight PoDM is $\Theta_{\kappa}\left(\log\kappa\right)$ and the tight PoA is $\Theta_{\kappa}\left((\log\kappa)^{2}\right)$.
@@
Under 6.1, if contribution distributions $F$ are MHR, the tight PoDM is in $[{\mathrm{e}},3.448]$ and the tight PoA is ${\mathrm{e}}^{2}$. Moreover, if $n=1$, the tight PoDM is ${\mathrm{e}}$.
@@ 7 Unknown Market Size
In practice, the principal may lack perfect knowledge of the number of agents or incur significant costs to acquire it. To address this, we introduce linear contract design for unknown market size, where the principal knows only a range for the number of agents, i.e., $n\in[s:\ell]$, rather than its exact value. Our goal is to determine how the principal in our three-party model should design linear contracts to safeguard her utility in this robust setting. To evaluate the principal’s utility under a given linear contract, we continue using the first-best and second-best benchmarks and redefine the price of anarchy and the price of double marginalization as follows:
@@
In the above definition, both ratios are defined in the worst case over the range $[s:\ell]$ of possible agent numbers. To ensure that these ratios have upper bounds related to the market size range, $[s:\ell]$, we propose two contract design schemes under regular and MHR contribution distributions, respectively. Our results show that when the market size is unknown, both ratios suffer a multiplicative loss compared to the bounds established for known market sizes in Section 5. Under regular distributions, the loss factor is $\frac{\ell}{s}$ (Theorem 7.1), whereas under MHR distributions, it reduces to $\frac{\log\ell}{\log s}$ (Theorem 7.2). Interestingly, the former bound is achieved by designing the contract assuming the largest possible market size, $\ell$, while the latter is obtained by assuming the smallest possible market size, $s$. Moreover, we provide an example illustrating that if the principal has no knowledge of the market size—i.e., if the number of agents ranges from $1$ to infinity—then, regardless of the contract design, the gap between her utility and these benchmarks cannot be upper bounded (Theorem 7.3).
@@
The intuition behind this theorem is that when $n\in[s:\ell]$, if the principal designs the contract as if the market size were $\ell$, then the worst-case upper bounds for the two ratios increase by a multiplicative factor of $\left\lceil\ell/s\right\rceil$ compared to the case where the principal knows the exact value of $n$. Moreover, when $s=\ell$, meaning the principal knows the exact market size, this result aligns with Theorem 5.1 and Theorem 5.6.
@@
Under MHR contribution distributions, Theorem 7.2 provides better results than Theorem 7.1. Meanwhile, when $s=\ell$, the result of this theorem is also consistent with the results of Theorem 5.6.
@@
For any small constant $\varepsilon>0$ and agents with identical expected rewards $r$ and MHR contribution distributions $F$, if the principal is only aware of the range of the number of agents $n\in[1:\infty]$, there exists no linear contract $\alpha$ such that both $\mathrm{PoA}(1,\infty)$ and $\mathrm{PoDM}(1,\infty)$ have upper bounds.
7.1 Analysis for Agents with a Regular Distribution
In this section, we prove Theorem 7.1. Our analysis relies on the following technical lemma, whose proof can be found in Section C.8.
Lemma 7.4.
For $n$ i.i.d. random variables $v_{i}~{}(i\in[n])$ distributed according to regular distribution $F$, if $n\in[s:\ell]$, we have
$\left\lceil\frac{\ell}{s}\right\rceil\cdot\left(1-F^{n}(z)\right)\geq\left(1-F^{\ell}(z)\right).$
Now we are ready to prove Theorem 7.1.
Proof of Theorem 7.1..
Consider the design of setting $\alpha^{{\dagger}}$ as the optimal contract for the number of agents being $\ell$, i.e.,
$\alpha^{{\dagger}}=\operatorname*{arg\,max}_{\alpha\in[0,1]}U\left(\alpha,n\right).$
We have
$z^{{\dagger}}=r\left(1-\alpha^{{\dagger}}\right)+\theta_{\alpha^{{\dagger}}}=\operatorname*{arg\,max}_{z\in\mathtt{supp}(F)}\phi(z)\left(1-F^{\ell}(z)\right).$
Thus, for any $n\in[s:\ell]$, we can further obtain that
$U_{\textsc{sb}}(n)\overset{(a)}{\leq}{}$ $U_{\textsc{sb}}(\ell)$
=$\overset{(b)}{=}{}$ $\mathcal{O}_{\kappa}\left(\log\kappa\right)\cdot\max\limits_{z\in\mathtt{supp}(F)}\phi(z)\left(1-F^{\ell}(z)\right)$
=$\overset{(c)}{=}{}$ $\mathcal{O}_{\kappa}\left(\log\kappa\right)\cdot\phi(z^{{\dagger}})\left(1-F^{\ell}(z^{{\dagger}})\right)$
=$\overset{(d)}{=}{}$ $\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left(\log\kappa\right)\cdot\phi(z^{{\dagger}})\left(1-F^{n}(z^{{\dagger}})\right)$
=$\overset{(e)}{=}{}$ $\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left(\log\kappa\right)\cdot U(\alpha^{{\dagger}},n),$
where inequality (a) holds since $U_{\textsc{fb}}(n)$ is non-decreasing as $n$ increases, equality (b) holds by Theorem 5.1, equality (c) holds due to the definition of $z^{{\dagger}}$, equality (d) holds by Lemma 7.4 and equality (e) holds by definition of $U\left(\alpha,n\right)$. Moreover, by Theorem 5.1, we can also obtain that
$U_{\textsc{fb}}(n)\leq U_{\textsc{fb}}(\ell)=\mathcal{O}_{\kappa}\left((\log\kappa)^{2}\right)\cdot\max\limits_{z\in\mathtt{supp}(F)}\phi(z)\left(1-F^{\ell}(z)\right)=\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left(\left(\log\kappa\right)^{2}\right)\cdot U(\alpha^{{\dagger}},n).$
Therefore, we can obtain that
$\mathrm{PoDM}(s,\ell)\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{sb}}(n)}{U(\alpha^{{\dagger}},n)}=\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left(\log\kappa\right),$
and
$\mathrm{PoA}(s,\ell)\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}=\frac{\ell}{s}\cdot\mathcal{O}_{\kappa}\left((\log\kappa)^{2}\right),$
which completes the proof. ∎
7.2 Analysis for Agents with an MHR Distribution
In this section, we prove Theorem 7.2. Our analysis relies on the following technical lemma, whose proofs can be found in Section C.9.
Lemma 7.5.
Now, we are ready to prove Theorem 7.2.
Proof of Theorem 7.2..
Suppose that for each agent $i\in[n]$, the contribution $w_{i}$ satisfies an MHR distribution $F$. According to Lemma 7.5, we have
For any $n\in[s:\ell]$, we finally obtain that
$\mathrm{PoA}(s,\ell)\leq\max_{n\in[s:\ell]}\frac{U_{\textsc{fb}}(n)}{U(\alpha^{{\dagger}},n)}\leq\frac{U_{\textsc{fb}}(\ell)}{U\left(\alpha^{{\dagger}},s\right)}\leq\frac{U_{\textsc{fb}}(\ell)}{U_{\textsc{fb}}(s)}\cdot\frac{U_{\textsc{fb}}(s)}{U\left(\alpha^{{\dagger}},s\right)}\overset{(a)}{\leq}\frac{\ln\left(\frac{1}{2}(1-F(0))\ell+1\right)}{\ln\left(\frac{1}{2}(1-F(0))s+1\right)}\cdot{\mathrm{e}}^{2}+o_{s,\ell}\left(1\right).$
where inequality (a) holds by the definition of $\alpha^{{\dagger}}$ and Theorem 5.6. Moreover, by Theorem 5.7, we can also obtain that
which completes the proof. ∎
@@ Appendix B Two Intermediaries with Different Mechanism Spaces
In our model, the principal’s utility depends critically on the constraints faced by the intermediary in her mechanism design spaces. A natural question arises: consider two intermediaries with different mechanism design spaces, if one’s mechanism space strictly contains the other’s, does the principal always benefit from the more flexible intermediary? We show that this is not universally true by offering two examples, illustrating that the inclusion-relation of mechanism design spaces for the two intermediaries does not imply the order of the principal’s optimal utility.
@@
Intermediary $A$ can choose all mechanisms while intermediary $B$ can choose a single mechanism that never assign the task to agents. In this example, principal has higher utility facing $A$ than $B$.
@@
Intermediary $A$ can choose all mechanisms while intermediary $B$ can choose a single mechanism: assigning the task to the agent with the lowest cost and paying her the second lowest cost among the other agents. In this example, principal has higher utility facing $B$ than $A$.
The intuition behind is that the utility of the intermediary depends both on the allocation and payment of the mechanism, but only the allocation directly affects the utility of the principal. Therefore, the intermediary that optimize her own utility within a smaller mechanism space may design allocation rules that are either more or less favorable to the principal.
@@ Appendix C Missing Proofs
C.1 Proof of Theorem 3.1
See 3.1
We first present how, when the contract is given by the principal, the mechanism between the intermediary and the $n$ agents can be converted into a single-item auction between a seller and $n$ buyers, and vice versa.
@@
Buyer $i$’s utility in the auction is given by $(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i})-(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-p_{\bm{t},i}(\bm{c}))=p_{\bm{t},i}(\bm{c})-c_{i}$ if agent $i$ is the winner in the mechanism, and 00 otherwise. This utility is identical to agent $i$’s utility in $\mathcal{M}_{\bm{t}}$. Additionally, since the mechanism $\mathcal{M}_{\bm{t}}$ and the auction $\mathcal{M}_{\bm{t}}^{{\dagger}}$ share the same winner, the truthfulness of $\mathcal{M}_{\bm{t}}$ implies the truthfulness of the auction $\mathcal{M}_{\bm{t}}^{{\dagger}}$, and vice versa. Then, we can observe that the seller’s revenue, given by the payment $\bm{t}\bm{\gamma}_{i^{\star}}-p_{\bm{t},i^{\star}}(\bm{c})$ from the winner $i^{\star}$, is equal to the utility of the intermediary in $\mathcal{M}_{\bm{t}}$.
@@
Note that the contracted contribution $w_{\bm{t},i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$ can be negative. Also, note that the contribution of agent $i$ is exactly her contracted contribution $w_{\bm{t},i}$ by setting $\bm{t}=\bm{1}$. We have assumed regularity of the contribution distribution in 2.4. Next, we show that this assumption implies the regularity of the contracted contribution distribution for any $\bm{t}$.
@@
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
@@
By Lemma C.2, the regularity property of the contribution distribution is preserved in the contracted contribution distribution. Since agents’ costs are independently drawn from the regular distributions $\bm{G}$, for a given contract $\bm{t}$, the corresponding values are also independently drawn from the regular distributions $\bm{F}_{\bm{t}}$. Therefore, by Lemma 2.1, the optimal mechanism for the intermediary is the virtual welfare maximizer mechanism. Then by Definition 2.3, we finish the proof. ∎
@@
See 4.1
@@
By the characterization in Theorem 3.1, the intermediary’s optimal mechanism is the virtual welfare maximizer mechanism with the value equal to the contracted contribution $w_{\alpha,i}$. According to the allocation rule of the virtual welfare maximizer mechanism, an agent is the winner if and only if her virtual value is non-negative and the highest among all agents. Therefore, the probability that the task is auctioned to agent $i$ is $\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}_{\bm{t}}\sim\bm{F}_{\bm{t}}}\!\left[{\phi_{\bm{t},i}(w_{\bm{t},i})\geq\max\nolimits_{j\in[n]}{\left(\phi_{\bm{t},j}(w_{\bm{t},j})\right)^{+}}}\right]$. Consequently, the principal’s problem of maximizing her utility in eq. 5 can be rewritten as
@@
Next, we consider the case where all agents have the identical expected reward profiles $\bm{\gamma}_{i}\equiv\bm{\gamma}$. Let $\bm{t}^{\star}$ be the optimal contract that achieves the maximum utility for the principal in eq. 5. Define $\alpha^{\star}\triangleq\frac{\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle}{r}$, and we have $\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle=\alpha^{\star}r$. Clearly, the optimality of $\bm{t}^{\star}$ ensures that $\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle\leq r$, which implies that $\alpha^{\star}\in[0,1]$. By Corollary 3.2, the principal’s utility remains unchanged if we replace the contract $\bm{t}^{\star}$ with the linear contract $\alpha^{\star}\cdot\bm{1}$.

### v2

@@ 5 Intermediary with Anonymous Pricing
Agents have identical expected reward profiles $\bm{\gamma}$ and contribution distributions $F$, and the mechanism space of the intermediary is restricted to anonymous pricing mechanisms.
@@
Under 5.1, we can obtain that given a contract $\bm{t}\in\mathbb{R}_{+}^{m}$, the intermediary’s expected utility under the anonymous price $p$ can be expressed as $\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\cdot\operatorname*{\mathrm{Pr}}\!\left[{\exists i\in[n],s.t.,c_{i}\leq p}\right]$. Note that the probability that there exists an agent with cost no more than the posted price $p$ is equal to the probability that there exists an agent with contribution no less than $\left\langle\bm{t},\bm{\gamma}\right\rangle-p$. The intermediary’s expected utility of an anonymous price $p$ is equal to $\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\cdot\left(1-F^{n}_{\bm{t}}\left(\left\langle\bm{t},\bm{\gamma}\right\rangle-p\right)\right)$. By definition of reserve price, the intermediary’s optimal anonymous price is $\left\langle\bm{t},\bm{\gamma}\right\rangle-\theta^{(n)}_{\bm{t}}$, where $\theta^{(n)}_{\bm{t}}$ is the reserve price of the distribution $F^{n}_{\bm{t}}$. Thus, the principal’s utility under contract $\bm{t}$ is equal to
$\left\langle\bm{1}-\bm{t},\bm{\gamma}\right\rangle\cdot\left(1-F^{n}_{\bm{t}}(\theta^{(n)}_{\bm{t}})\right)$.
Similar to results in Section 3, we show that when agents are ex ante symmetric, the optimal contract is linear and its induced payoff for the principal can be considered as a variant of virtual value pricing (Theorem 5.1). The proof is similar to the one for Theorem 3.1, and is deferred to Section C.6.
@@
Under 5.1, there exists an optimal contract that is linear. In particular, the optimal payoff of the principal (denoted by $U_{\mathrm{posted}}$) is the optimal objective value of the optimization program defined as follows:
@@
Under 5.1, Theorem 5.1 implies that the principal’s problem of designing an optimal contract remains a virtual value pricing problem as established Part (iii) of Theorem 3.1. The only difference is that the corresponding virtual value function becomes that of the distribution $F^{n}$. Intuitively, this virtual value pricing can be viewed as posting a price to a single virtual agent whose value is equal to the first-order statistic among those of the real agents. If she buys at that price, a payment equal to her virtual value will be received. Note that the principal’s utility in this case is no more than that in the case where there is no restriction on the intermediary’s mechanism space. A natural question is whether there is any monotonicity in the principal’s utility regarding two intermediaries with an inclusion relationship between their mechanism spaces? We offer two negative examples for this question in Appendix B.
We next analyze the impact of the intermediary measured by PoA and PoDM as we have done in Section 4.161616Here we define PoA and PoDM under 5.1 by $\mathrm{PoA}\triangleq\frac{U_{\textsc{fb}}}{U_{\mathrm{posted}}}$ and $\mathrm{PoDM}\triangleq\frac{U_{\textsc{sb}}}{U_{\mathrm{posted}}}$, where benchmarks $U_{\textsc{fb}},U_{\textsc{sb}}$ follow the same definitions in Section 4. Interestingly, as shown in Theorems 5.2 and 5.3, even when the intermediary is restricted to the space of anonymous pricing mechanisms, the (worst-case) PoDM and PoA are the same as the setting where the intermediary can design any DSIC, ex-post IR mechanisms. The proofs of both theorems are similar to their analogy in Section 4 and thus are deferred into Sections C.7 and C.8, respectively.
@@
Under 5.1, if contribution distributions $F$ are regular with support $\mathtt{supp}(F)=\left[\underline{w},\bar{w}\right]$ $\left(0<\underline{w}\leq\bar{w}\leq r\right)$, if $1\leq\bar{w}/\underline{w}\leq\kappa$, the tight PoDM is $\Theta_{\kappa}\left(\log\kappa\right)$ and the tight PoA is $\Theta_{\kappa}\left((\log\kappa)^{2}\right)$.
@@
Under 5.1, if contribution distributions $F$ are MHR, the tight PoDM is in ${\mathrm{e}},3.448]$ and the tight PoA is ${\mathrm{e}}^{2}$. Moreover, if $n=1$, the tight PoDM is $\mathrm{e}$}.
@@ 6 Unknown Market Size
In practice, the principal may lack perfect knowledge of the number of agents or incur high costs to acquire it. Motivated by this, we introduce and explore the linear contract design for unknown market size, where the principal knows only a range for the number of agents, i.e., $n\in[s:\ell]$, rather than its exact value. Our goal is to determine how the principal in our three-party model should design linear contracts to safeguard her utility in this robust setting. To evaluate the principal’s utility under a given linear contract, we continue using the first-best and second-best benchmarks and redefine the price of anarchy and the price of double marginalization as follows:
@@
In the above definition, both ratios are defined in the worst case over the range $s:\ell]$ of possible agent numbers. To ensure that these ratios have upper bounds related to the market size range, $s:\ell]$, we propose two contract design schemes under regular and MHR contribution distributions, respectively. Our results show that when the market size is unknown, both ratios suffer a multiplicative loss compared to the bounds established for known market sizes in Section 4. Under regular distributions, the loss factor is $\frac{\ell}{s}$ (Theorem 6.1), whereas under MHR distributions, it reduces to $\frac{\log\ell}{\log s}$ (Theorem 6.2). Notably, the former bound arises from optimizing the contract for any market size in $s:\ell]$, while the latter is uniquely attained by optimizing for the minimal size ss—optimizing for any other size fails to provide this guarantee. Moreover, we provide an example illustrating that if the principal has no knowledge of the market size—i.e., if the number of agents ranges from $1$ to infinity—then, regardless of the contract design, the gap between her utility and these benchmarks cannot be upper bounded (Theorem 6.3).
@@
The intuition behind this theorem is that when $n\in[s:\ell]$, if the principal designs the contract as if the market size were $\ell$, then the worst-case upper bounds for the two ratios increase by a multiplicative factor of $\left\lceil\ell/s\right\rceil$ compared to the case where the principal knows the exact value of $n$. Moreover, when $s=\ell$, meaning the principal knows the exact market size, this result aligns with Theorem 4.1 and Theorem 4.2. The detailed proof of Theorem 6.1 can be found in Section C.9.
@@
Under MHR contribution distributions, Theorem 6.2 provides better results than Theorem 6.1. Meanwhile, when $s=\ell$, the result of this theorem is also consistent with the results of Theorem 4.2. The detailed proof of Theorem 6.2 can be found in Section C.10.
@@
For any small constant $\varepsilon>0$ and agents with identical expected rewards $r$ and MHR contribution distributions $F$, if the principal is only aware of the range of the number of agents $n\in[1:\infty]$, no linear contract $\alpha$ satisfies that both $\mathrm{PoA}(1,\infty)$ and $\mathrm{PoDM}(1,\infty)$ have upper bounds.
@@ Appendix B Two Intermediaries with Different Mechanism Spaces
In our model, the principal’s utility depends critically on the constraints faced by the intermediary in her mechanism design spaces. A natural question arises: consider two intermediaries with different mechanism design spaces. If one’s mechanism space strictly contains the other’s, does the principal always benefit from the more flexible intermediary? We show that this is not universally true by offering two examples, illustrating that the inclusion-relation of mechanism design spaces for the two intermediaries does not imply the order of the principal’s optimal utility.
@@
Intermediary $A$ can choose all mechanisms while intermediary $B$ can choose a single mechanism that never assigns the task to agents. In this example, the principal has higher utility facing $A$ than $B$.
@@
Intermediary $A$ can choose all mechanisms while intermediary $B$ can choose a single mechanism: assigning the task to the agent with the lowest cost and paying her the second lowest cost among the other agents. In this example, the principal has higher utility facing $B$ than $A$.
The intuition behind this is that the utility of the intermediary depends both on the allocation and payment of the mechanism, but only the allocation directly affects the utility of the principal. Therefore, the intermediary that optimize her own utility within a smaller mechanism space may design allocation rules that are either more or less favorable to the principal.
@@ Appendix C Missing Proofs
C.1 Proof of Proposition 3.2
In this section, we will prove Proposition 3.2.
See 3.2
An immediate implication from the above result is as follows.
Corollary C.1.
At a high level, to prove Proposition 3.2, we first show that any truthful mechanism between the intermediary and $n$ agents can be converted into a truthful single-item auction between a seller and $n$ buyers, and vice versa, as described in Lemma C.3. The key is to convert agent $i$’s private cost $c_{i}$ into buyer $i$’s private value $v_{i}$ in the auction. The contract $\bm{t}$ plays a pivotal role in this conversion: $v_{i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$, which is the contracted contribution of agent $i$. Thus, the seller’s revenue in the auction meets the intermediary’s utility in the mechanism. Then, we establish that the regularity or MHR property of the contribution distribution is preserved in the distribution of the contracted contribution for any contract $\bm{t}$, as shown in Lemma C.4. This crucial property enables us to apply standard auction theory, demonstrating that the optimal mechanism for the intermediary is indeed the virtual welfare maximizer mechanism as follows.
Consider the multi-buyer scenario where a seller wants to sell a single item to one of $n$ buyers, whose private values of the item, $\left\{v_{i}\right\}_{i\in[n]}$, are drawn independently from regular distributions $\left\{F_{i}\right\}_{i\in[n]}$. Myerson (1981) proved that the virtual welfare maximizer mechanism is DSIC, ex-post IR, and maximizes the seller’s expected revenue.
Definition C.1 (Virtual welfare maximizer, Myerson, 1981).
  1. 1.
  2. 2.
  3. 3.
    $p_{i}^{{\dagger}}(\bm{v})\triangleq\phi_{i}^{-1}\left(\max\nolimits_{j\in[n]\setminus\left\{i\right\}}{\left(\phi_{j}\left(v_{j}\right)\right)^{+}}\right)\cdot x^{{\dagger}}_{i}(\bm{v}),$
Lemma C.2 (Myerson, 1981).
To prove Proposition 3.2, we first present how, when the contract is given by the principal, the mechanism between the intermediary and the $n$ agents can be converted into a single-item auction between a seller and $n$ buyers, and vice versa.
@@
Buyer $i$’s utility in the auction is given by $(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i})-(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-p_{\bm{t},i}(\bm{c}))=p_{\bm{t},i}(\bm{c})-c_{i}$ if agent $i$ is the winner in the mechanism, and 0 otherwise. This utility is identical to agent $i$’s utility in $\mathcal{M}_{\bm{t}}$. Additionally, since the mechanism $\mathcal{M}_{\bm{t}}$ and the auction $\mathcal{M}_{\bm{t}}^{{\dagger}}$ share the same winner, the truthfulness of $\mathcal{M}_{\bm{t}}$ implies the truthfulness of the auction $\mathcal{M}_{\bm{t}}^{{\dagger}}$, and vice versa. Then, we can observe that the seller’s revenue, given by the payment $\bm{t}\bm{\gamma}_{i^{\star}}-p_{\bm{t},i^{\star}}(\bm{c})$ from the winner $i^{\star}$, is equal to the utility of the intermediary in $\mathcal{M}_{\bm{t}}$.
@@
Note that the contracted contribution $w_{\bm{t},i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$ can be negative. Also, note that the contribution of agent $i$ is exactly her contracted contribution $w_{\bm{t},i}$ by setting $\bm{t}=\bm{1}$. We have assumed regularity of the contribution distribution in 2.3. Next, we show that this assumption implies the regularity of the contracted contribution distribution for any $\bm{t}$.
@@
Now we are ready to prove Proposition 3.2.
Proof of Proposition 3.2.
@@
By Lemma C.4, the regularity property of the contribution distribution is preserved in the contracted contribution distribution. Since agents’ costs are independently drawn from the regular distributions $\bm{G}$, for a given contract $\bm{t}$, the corresponding values are also independently drawn from the regular distributions $\bm{F}_{\bm{t}}$. Therefore, by Lemma C.2, the optimal mechanism for the intermediary is the virtual welfare maximizer mechanism. Then by Definition C.1, we finish the proof. ∎
@@
In this section, we will prove Theorem 3.1.
See 3.1
To prove Theorem 3.1, we first analyze the continuity of the monopoly reserve as follows.
Lemma C.5 (Continuity of monopoly reserve).
For any fixed agent $i\in[n]$, consider the monopoly reserve price $\theta_{\alpha,i}$ as a function of the linear contract $\alpha$. Then, $\theta_{\alpha,i}$ is continuous in w.r.t., $\alpha\in[0,1]$.
@@
By the definition of reserve price, we have
Define two functions $\upsilon(\alpha)\triangleq-\theta_{\alpha}+\alpha\bm{\gamma}$ and $\varphi(z)\triangleq z+\frac{G(z)}{G^{\prime}(z)}$. By replacing $\theta_{\alpha}$ in eq. 1 with $\upsilon(\alpha)+\alpha\bm{\gamma}$, we get that for any $\alpha\in[0,1]$, it holds
Clearly, $\varphi(\upsilon(\cdot))$ is continuous in $[0,1]$. By 2.3, $\varphi(\cdot)$ is continuous in $\mathtt{supp}(G)$. Hence, we have $\upsilon(\cdot)$ is continuous in $[0,1]$. Since $\theta_{\alpha}=-\upsilon(\alpha)+\alpha\bm{\gamma}$, we finally obtain that $\theta_{\alpha}$ is also continuous in $[0,1]$. This completes our proof. ∎
Now, we are ready to prove Theorem 3.1.
Proof of Theorem 3.1..
@@
By the characterization in Proposition 3.2, the intermediary’s optimal mechanism is the virtual welfare maximizer mechanism with the value equal to the contracted contribution $w_{\alpha,i}$. According to the allocation rule of the virtual welfare maximizer mechanism, an agent is the winner if and only if her virtual value is non-negative and the highest among all agents. Therefore, the probability that the task is auctioned to agent $i$ is $\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}_{\bm{t}}\sim\bm{F}_{\bm{t}}}\!\left[{\phi_{\bm{t},i}(w_{\bm{t},i})\geq\max\nolimits_{j\in[n]}{\left(\phi_{\bm{t},j}(w_{\bm{t},j})\right)^{+}}}\right]$. Consequently, the principal’s problem of maximizing her utility in eq. 2 can be rewritten as
@@
Next, we consider the case where all agents have the identical expected reward profiles $\bm{\gamma}_{i}\equiv\bm{\gamma}$. Let $\bm{t}^{\star}$ be the optimal contract that achieves the maximum utility for the principal in eq. 2. Define $\alpha^{\star}\triangleq\frac{\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle}{r}$, and we have $\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle=\alpha^{\star}r$. Clearly, the optimality of $\bm{t}^{\star}$ ensures that $\left\langle\bm{t}^{\star},\bm{\gamma}\right\rangle\leq r$, which implies that $\alpha^{\star}\in[0,1]$. By Corollary C.1, the principal’s utility remains unchanged if we replace the contract $\bm{t}^{\star}$ with the linear contract $\alpha^{\star}\cdot\bm{1}$.